# Mapping Dark Matter Through the Dust of the Milky Way Part I: Dust Correction and Phase Space Density

### v1

@@ 1 Introduction
Dark matter is unequivocal evidence of physics beyond the Standard Model. Though the nature of dark matter is of fundamental interest to particle physics, it must be remembered that all lines of evidence for the existence of dark matter – as well as nontrivial constraints on its particle interactions – come from astrophysics and cosmology [1, 2, 3, 4, 5, 6, 7]. Of particular relevance to this work, the density profile of the Milky Way’s dark matter halo is an important component of studies of dark matter. For example, the local density of our dark matter halo sets the expected scattering rate with visible matter in dark matter direct detection experiments (see Ref. [8] for a compilation of recent measurements and Refs. [9, 10] for detailed discussions). Additionally, non-trivial modifications of the dark sector, such as self-interactions or a component of “warm” dark matter, can alter the expected halo density profile towards the Galactic center (see Refs. [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]).
@@
Historically, measurements of $f$ (and its derivatives) with sufficient accuracy to yield stable and tractable solutions to the CBE were not possible, due in large part to the high dimensionality (six) of the phase space. Instead, moments of the velocity are obtained from Eq. (1.1) under the assumption of equilibrium ($\partial f/\partial t=0$). The first such moments result in the Jeans Equations. These equations, along with binning stellar kinematic data under assumptions of symmetry, have formed the basis of many of the measurements of the Milky Way potential [23, 24, 25, 26, 27, 28].
Machine learning algorithms, in particular normalizing flows (see Refs. [29, 30]), allow modeling of $f$ (and its gradients) from sparser data than previously possible. This opens a new avenue to solving the CBE for $\Phi$ directly, using flows trained on data to represent the previously-inaccessible PSD. Additionally, flow models of $f$ are flexible and do not require any symmetry assumptions, allowing for a data-driven measurement of $\Phi$ with minimal assumptions.
@@
Taking the second derivative of $\Phi$ results in the mass density $\rho$ within the Galaxy via the Poisson Equation. This technique was proposed and applied to mock potentials and stellar distributions in Refs. [31, 32, 33], and demonstrated in Ref. [34] using synthetic stars drawn from a fully cosmological $N$-body simulation with realistic errors. In Ref. [8], the authors of this work applied this technique to Gaia’s $3^{\rm rd}$ Data Release (DR3) [35, 36]. In this measurement, normalizing flows were trained on the six-dimensional kinematics of nearby stars as measured by the Gaia Space Observatory’s [37] radial velocity spectrometer (RVS) to provide the estimates of $f(\vec{x},\vec{v})$. Using this technique, we presented a data-driven map of the dark matter density in select regions of the local Milky Way. However, this study was limited by several factors:
@@
    The acceleration $\vec{a}$ was computed pointwise across the analysis volume instead of obtaining a continuous function. Consequently, we could not enforce the curl-free condition ×$\nabla\times\vec{a}=0$, nor could we prioritize solutions with a positive-definite mass density $\rho>0$.
@@
Figure 1: Projection of Gaia DR3 RC/RGB stars within 4 kpc used for training the normalizing flows used in this analysis. The names and locations of prominent nearby dust clouds (dark nebulae (DNe) and molecular clouds (MoC)) are overlaid to contextualize the irregular and dark low-number density patches of the sky. The dark patch at Baade’s window near the Parrot’s Head Nebula $(\ell,b)\approx(1^{\circ},-3.66^{\circ})$ is an extremely crowded star field near the Galactic center. This crowding saturates Gaia’s instruments, decreasing observational completeness in this region (Ref. [38]).
@@
This suppression of the observed PSD $f_{\rm obs}$ by dust makes it difficult to use Boltzmann techniques to measure the gravitational acceleration and potential (as well as the dark matter mass density) in the disk and towards the Galactic center. In Ref. [8], we restricted ourselves to measurements along lines of sight either perpendicular to the disk or offset from it, in order to avoid the most significant areas of dust.
The dusty features of the Milky Way and their implications for the completeness of surveys such as Gaia have been well-studied (Refs. [39, 40, 41, 42]). However, using these dust maps to infer the true PSD of stars from the observed values is difficult. Inverting an extinction function to create a selection efficiency function requires knowledge of the concealed stellar population. Additionally, small errors in estimating this observational selection function or efficiency will result in large errors in the derivatives of the PSD. These errors then propagate in to the inferred values of the gravitational acceleration through the CBE.
@@
$\vec{v}\cdot\vec{\nabla}\ln f_{\rm obs}-\vec{v}\cdot\vec{\nabla}\ln\epsilon(\vec{x})+\vec{\nabla}\Phi(\vec{x})\cdot\frac{\partial\ln f_{\rm obs}}{\partial\vec{v}}=0.$ (1.3)
@@
This novel approach to dust correction addresses point 1 above and extends our previous work to nearly the entire $4$ kpc ball around the Solar location. To address points 2 and 3 above, we choose to parametrize both the log-efficiency $\ln\epsilon$ and the gravitational potential $\Phi$ with feed-forward neural networks (NN). This automatically guarantees the curl-free condition for $\vec{a}(\vec{x})$ and also allows us to include physical constraints such as positive-definite mass density. In addition, the use of a free-form model for $\Phi$ balances the computational efficiency of a global model for $\Phi$ (as opposed to the pointwise approach adopted in our previous work) without introducing biases that arise from restrictions to smaller spaces of functional forms. The idea of parameterizing $\Phi$ with a NN was first proposed in Ref. [31], but this is the first work to introduce an efficiency function into the CBE and simultaneously learn both $\Phi$ and $\epsilon$.
The outline of our paper is as follows. In Section 2, we describe the dataset used to learn the observed six-dimensional PSD $f_{\rm obs}$ of the local Milky Way. In Section 3, we review and describe updates to the masked autoregressive flow (MAF) architecture and training techniques introduced in our prior work Ref. [8], and then describe the training regimen for the $\Phi(\vec{x})$ and $\epsilon(\vec{x})$ neural networks. In Section 4, we compare our learned $\epsilon(\vec{x})$ to a state-of-the-art three-dimensional dust map, and then examine the dust-corrected phase space density. We reserve a detailed analysis of $\Phi(\vec{x})$ and downstream observables such as the three-dimensional acceleration and dark matter density fields for a companion paper, Ref. [43].
@@ 2.1 Coordinate System
Our Cartesian coordinate system is a right-handed Galactocentric coordinate system, where the center of the Galaxy is at the origin. The Sun lies close to the $x$ axis at $x=8.122$ kpc and $y=0$ kpc. The Galactic midplane is defined as $z=0$, with the Sun being slightly off-set above the disk at $z=0.0208$ kpc. The Large and Small Magellanic Clouds are below the disk, with $z<0$. At the Sun’s location, the positive $y$ axis is oriented counter to the direction of rotation of the disk. [^1] This is equivalent to a $180^{\circ}$ rotation of the traditional Galactocentric coordinates used in Astropy [44, 45].
@@ 2.2 Selecting a Complete RC/RGB Dataset
We use the same dataset of high-luminosity stars within 4 kpc of the Solar location that we considered previously in Ref. [8]; this consists primarily of Red Clump (RC) and Red Giant Branch (RGB) stars. The RC is of particular interest in this work, as well as other recent studies (Refs. [24, 46]) of equilibrium Galactic dynamics, since it is expected to be composed of relatively old stars. These old stars have had sufficient time to reach equilibrium, in the absence of any recent external dynamic perturbations such as mergers [9].
@@
This criterion imposes an explicit cut on absolute magnitude $M_{G}<0.990$. The location of this cut with respect to prominent stellar populations in the Gaia color-magnitude space is visualized in Figure 2. Prior to the selection on $M_{G}$, there are 24,789,061 well-measured stars in the Gaia DR3 dataset within 4 kpc of the Sun which have full six-dimensional kinematics. This number is reduced to 5,811,956 by the magnitude selection, a majority of which ($\sim 69\%$) lie within the RC. The diagonal smearing of the color-magnitude diagram in Figure 2 is caused by dust extinction. The vertical component of the smearing is due to the overall dimming of stars in the RVS band (increasing $M_{G}$), whereas the horizontal component of the smearing is due to “reddening”, the relative difference in dust extinction between the red and blue photometric bands of Gaia (redder is increasing $BP-RP$).
We intentionally do not extinction-correct the primary dataset used in this work, as these corrections can only be applied to dimmed stars that are nearby enough for Gaia to observe. Instead, we outline a novel technique for post-hoc dust correction in the next section. In Section 4, we compare the machine-learned efficiency factor with the expectation from previous dust map studies. In particular, we use the dust map found in Ref. [42] to correct the $M_{G}$ of nearby stars to develop an independent expectation for the missing fraction of stars in nearby regions of space. We refer to this dust map as “L22” for the remainder of this work, and describe the construction of the efficiency function from L22 in Appendix B.
@@ 3.1 MAF Models
We train Masked Autoregressive Flows (MAFs, Ref. [47]) to learn the components of $f_{\rm obs}$ as well as their gradients:
@@
where $n(\vec{x})$ is the number density and $p(\vec{v}|\vec{x})$ is the conditional velocity distribution of the RC/RGB dataset. MAFs, a type of normalizing flow, model the density of data $p(x)$ by learning an invertible and differentiable transformation from a simple latent base distribution to $p(x)$. They are constructed from a composition of transformations that satisfy the autoregressive property, enabling a tractable Jacobian for the transformation. This property allows us to efficiently compute gradients of $f_{\rm obs}$ with respect to $(\vec{x},\vec{v})$ by differentiating the MAFs directly with respect to their input parameters. For a detailed discussion of normalizing flows, see Ref. [29, 30].
Figure 2: Absolute magnitude $M_{G}$ versus color $BP-RP$ of all Gaia DR3 stars within 4 kpc of the Sun with full kinematic information, photometric magnitudes, and relative parallax errors $\delta\varpi/\varpi<0.33$. The 4 kpc completeness cut is shown as a white dashed line: stars above this line outside the hatched region are used in our MAF training sample. The Red Clump (RC) – the largest population used in this analysis – is outlined in a white box. Figure reproduced from the authors’ previous work, Ref. [8].
The MAFs used in this work are similar, with minor updates, to those used in our previous work (Ref. [8]). The modifications are:
@@
    Our data preprocessing has been updated to eliminate bias from the autoregressive structure of the MAF. As noted in Ref. [8], MAFs are sensitive to the provided ordering of inputs (in our case, the ordering of Cartesian directions). To counteract this, we inserted Cartesian coordinate permutations between the layers of the MAF. However, since these permutations were identical across training realizations, it is possible that some residual autoregressive bias remained. In this work, we have introduced a further random rotation of the input $\{\vec{x},\vec{v}\}$ of the RC/RGB dataset after centering (${\overline{x}_{i}}^{\prime}=0$) and standardization ($\sigma_{x_{i}}^{\prime}=1$) for each training realization.
@@
Aside from these minor changes, the other details of the MAF models are unchanged, in particular they have the same size and structure as in Ref. [8]. The MAFs for both $n(\vec{x})$ and $p(\vec{v}|\vec{x})$ are composed of two MADE blocks, each with 10 hidden layers that are 48 nodes wide. All MAF models are trained by minimizing the negative log-likelihood (NLL) of the Gaia DR3 RC/RGB dataset introduced in Section 2. For a complete discussion of the MAF training procedure, refer to Ref. [8].
@@ 3.2 Training Dust Efficiency and Potential
We parameterize the functions $\Phi(\vec{x})$ and $\ln\epsilon(\vec{x})$ with fully-connected NNs. Both use five hidden layers, each 100 neurons wide, with the same ${\rm GINT}$ activations used for the MAFs. An additional transformation layer is included in the $\ln\epsilon$ network to smoothly map the output from $\mathbb{R}$ to $(-\infty,0)$ such that $\epsilon$ is constrained within $(0,1)$. [^2] We choose the $C^{\infty}$ function $y=-\frac{1}{\pi}\left(x\left(\arctan{x}+\pi/2\right)+1\right)$ for this purpose. The two networks are trained by minimizing the mean-square error (MSE) of CBE in Eq. (1.3) with respect to the NN parameters $\bm{\theta}$ and $\bm{\vartheta}$:

### v2

# Sweeping the Dust Away – Correcting the Phase Space Density of the Milky Way with Unsupervised Machine Learning

@@ 1 Introduction
Dark matter is unequivocal evidence of physics beyond the Standard Model. Though the nature of dark matter is of fundamental interest to particle physics, it must be remembered that all lines of evidence for the existence of dark matter – as well as nontrivial constraints on its particle interactions – come from astrophysics and cosmology [1980ApJ...238..471R, 1939LicOB..19...41B, Salucci:2018hqu, 2011ARA&A..49..409A, 1933AcHPh...6..110Z, Planck:2018vyg, Clowe:2003tk]. Of particular relevance to this work, the density profile of the Milky Way’s dark matter halo is an important component of studies of dark matter. For example, the local density of our dark matter halo sets the expected scattering rate with visible matter in dark matter direct detection experiments (see Ref. [lim2023mapping] for a compilation of recent measurements and Refs. [2014JPhG...41f3101R, 2017MNRAS.465...76M] for detailed discussions). Additionally, non-trivial modifications of the dark sector, such as self-interactions or a component of “warm” dark matter, can alter the expected halo density profile towards the Galactic center (see Refs. [Buckley:2017ijx, 2000ApJ...534L.143B, 2000ApJ...543..514K, 2000ApJ...544L..87Y, 2001ApJ...547..574D, 2002ApJ...581.. [^7] C, 2011MNRAS.415.1125K, 2012MNRAS.423.3740V, 2013MNRAS.430...81R, 2013MNRAS.430..105P, 2013MNRAS.431L..20Z, 2018ApJ...853..109E]).
@@
Historically, measurements of $f$ (and its derivatives) with sufficient accuracy to yield stable and tractable solutions to the CBE were not possible, due in large part to the high dimensionality (six) of the phase space. Instead, moments of the velocity are obtained from Eq. (1.1) under the assumption of equilibrium ($\partial f/\partial t=0$). The first such moments result in the Jeans Equations. These equations, along with binning stellar kinematic data under assumptions of symmetry, have formed the basis of many of the measurements of the Milky Way potential [2018MNRAS.478.1677S, 2020A&A...643A..75S, 2020MNRAS.494.6001N, 2021ApJ...916..112N, 2020MNRAS.495.4828G, 2018A&A...615A..99H].
Machine learning algorithms, in particular normalizing flows (see Refs. [9089305, 2019arXiv191202762P]), allow modeling of $f$ (and its gradients) from sparser data than previously possible. This opens a new avenue to solving the CBE for $\Phi$ directly, using flows trained on data to represent the previously-inaccessible PSD. Additionally, flow models of $f$ are flexible and do not require any symmetry assumptions, allowing for a data-driven measurement of $\Phi$ with minimal assumptions.
@@
Taking the second derivative of $\Phi$ results in the mass density $\rho$ within the Galaxy via the Poisson Equation. This technique was proposed and applied to mock potentials and stellar distributions in Refs. [2023ApJ...942...26G, 2021MNRAS.506.5721A, 10.1093/mnras/stac153], and demonstrated in Ref. [2023MNRAS.521.5100B] using synthetic stars drawn from a fully cosmological $N$-body simulation with realistic errors. In Ref. [lim2023mapping], the authors of this work applied this technique to Gaia’s $3^{\rm rd}$ Data Release (DR3) [2021Gaia, 2022arXiv220800211G]. In this measurement, normalizing flows were trained on the six-dimensional kinematics of nearby stars as measured by the Gaia Space Observatory’s [2016Gaia] radial velocity spectrometer (RVS) to provide the estimates of $f(\vec{x},\vec{v})$. Using this technique, we presented a data-driven map of the dark matter density in select regions of the local Milky Way. However, this study was limited by several factors:
@@
    The acceleration $\vec{a}$ was computed pointwise across the analysis volume instead of obtaining a continuous function. Consequently, we could not enforce the curl-free condition $\nabla\times\vec{a}=0$, nor could we prioritize solutions with a positive-definite mass density $\rho>0$.
@@
Figure 1: Projection of Gaia DR3 RC/RGB stars within 4 kpc used for training the normalizing flows used in this analysis. The names and locations of prominent nearby dust clouds (dark nebulae (DNe) and molecular clouds (MoC)) are overlaid to contextualize the irregular and dark low-number density patches of the sky. The dark patch at Baade’s window near the Parrot’s Head Nebula $(\ell,b)\approx(1^{\circ},-3.66^{\circ})$ is an extremely crowded star field near the Galactic center. This crowding saturates Gaia’s instruments, decreasing observational completeness in this region (Ref. [2021A&A...649A...5F]).
@@
This suppression of the observed PSD $f_{\rm obs}$ by dust makes it difficult to use Boltzmann techniques to measure the gravitational acceleration and potential (as well as the dark matter mass density) in the disk and towards the Galactic center. In Ref. [lim2023mapping], we restricted ourselves to measurements along lines of sight either perpendicular to the disk or offset from it, in order to avoid the most significant areas of dust.
The dusty features of the Milky Way and their implications for the completeness of surveys such as Gaia have been well-studied (Refs. [2023A&A...674A..31D, 2023A&A...669A..55C, 2020PASP..132g4501R, Gaia2mass2022]). However, using these dust maps to infer the true PSD of stars from the observed values is difficult. Inverting an extinction function to create a selection efficiency function requires knowledge of the concealed stellar population. Additionally, small errors in estimating this observational selection function or efficiency will result in large errors in the derivatives of the PSD. These errors then propagate in to the inferred values of the gravitational acceleration through the CBE.
@@
$\vec{v}\cdot\vec{\nabla}\ln f_{\rm obs}-\vec{v}\cdot\vec{\nabla}\ln\epsilon(\vec{x})-\vec{\nabla}\Phi(\vec{x})\cdot\frac{\partial\ln f_{\rm obs}}{\partial\vec{v}}=0.$ (1.3)
@@
This novel approach to dust correction addresses point 1 above and extends our previous work to nearly the entire $4$ kpc ball around the Solar location. To address points 2 and 3 above, we choose to parametrize both the log-efficiency $\ln\epsilon$ and the gravitational potential $\Phi$ with feed-forward neural networks (NN). This automatically guarantees the curl-free condition for $\vec{a}(\vec{x})$ and also allows us to include physical constraints such as positive-definite mass density. In addition, the use of a free-form model for $\Phi$ balances the computational efficiency of a global model for $\Phi$ (as opposed to the pointwise approach adopted in our previous work) without introducing biases that arise from restrictions to smaller spaces of functional forms. The idea of parameterizing $\Phi$ with a NN was first proposed in Ref. [2023ApJ...942...26G], but this is the first work to introduce an efficiency function into the CBE and simultaneously learn both $\Phi$ and $\epsilon$.
The outline of our paper is as follows. In Section 2, we describe the dataset used to learn the observed six-dimensional PSD $f_{\rm obs}$ of the local Milky Way. In Section 3, we review and describe updates to the masked autoregressive flow (MAF) architecture and training techniques introduced in our prior work Ref. [lim2023mapping], and then describe the training regimen for the $\Phi(\vec{x})$ and $\epsilon(\vec{x})$ neural networks. In Section 4, we compare our learned $\epsilon(\vec{x})$ to a state-of-the-art three-dimensional dust map, and then examine the dust-corrected phase space density. We reserve a detailed analysis of $\Phi(\vec{x})$ and downstream observables such as the three-dimensional acceleration and dark matter density fields for a companion paper, Ref. [dustpaperII].
@@ 2.1 Coordinate System
Our Cartesian coordinate system is a right-handed Galactocentric coordinate system, where the center of the Galaxy is at the origin. The Sun lies close to the $x$ axis at $x=8.122$ kpc and $y=0$ kpc. The Galactic midplane is defined as $z=0$, with the Sun being slightly off-set above the disk at $z=0.0208$ kpc. The Large and Small Magellanic Clouds are below the disk, with $z<0$. At the Sun’s location, the positive $y$ axis is oriented counter to the direction of rotation of the disk. [^1] This is equivalent to a $180^{\circ}$ rotation of the traditional Galactocentric coordinates used in Astropy [astropy:2013, astropy:2018].
@@ 2.2 Selecting a Complete RC/RGB Dataset
We use the same dataset of high-luminosity stars within 4 kpc of the Solar location that we considered previously in Ref. [lim2023mapping]; this consists primarily of Red Clump (RC) and Red Giant Branch (RGB) stars. The RC is of particular interest in this work, as well as other recent studies (Refs. [2020A&A...643A..75S, 2024arXiv240608158B]) of equilibrium Galactic dynamics, since it is expected to be composed of relatively old stars. These old stars have had sufficient time to reach equilibrium, in the absence of any recent external dynamic perturbations such as mergers [2014JPhG...41f3101R].
@@
This criterion imposes an explicit cut on absolute magnitude $M_{G}<0.990$. The location of this cut with respect to prominent stellar populations in the Gaia color-magnitude space is visualized in Figure 2. Prior to the selection on $M_{G}$, there are 24,789,061 well-measured stars in the Gaia DR3 dataset within 4 kpc of the Sun which have full six-dimensional kinematics. This number is reduced to 5,811,956 by the magnitude selection, a majority of which ($\sim 69\%$) lie within the RC. The diagonal smearing of the color-magnitude diagram in Figure 2 is caused by dust extinction. The vertical component of the smearing is due to the overall dimming of stars in the RVS band (increasing $M_{G}$), whereas the horizontal component of the smearing is due to “reddening”, the relative difference in dust extinction between the red and blue photometric bands of Gaia (redder is increasing BPRPBP-RP).
We intentionally do not extinction-correct the primary dataset used in this work, as these corrections can only be applied to dimmed stars that are nearby enough for Gaia to observe. Instead, we outline a novel technique for post-hoc dust correction in the next section. In Section 4, we compare the machine-learned efficiency factor with the expectation from previous dust map studies. In particular, we use the dust map found in Ref. [Gaia2mass2022] to correct the $M_{G}$ of nearby stars to develop an independent expectation for the missing fraction of stars in nearby regions of space. We refer to this dust map as “L22” for the remainder of this work, and describe the construction of the efficiency function from L22 in Appendix B.
@@ 3.1 MAF Models
We train Masked Autoregressive Flows (MAFs, Ref. [papamakarios2018masked]) to learn the components of $f_{\rm obs}$ as well as their gradients:
@@
where $n(\vec{x})$ is the number density and $p(\vec{v}|\vec{x})$ is the conditional velocity distribution of the RC/RGB dataset. MAFs, a type of normalizing flow, model the density of data $p(x)$ by learning an invertible and differentiable transformation from a simple latent base distribution to $p(x)$. They are constructed from a composition of transformations that satisfy the autoregressive property, enabling a tractable Jacobian for the transformation. This property allows us to efficiently compute gradients of $f_{\rm obs}$ with respect to $(\vec{x},\vec{v})$ by differentiating the MAFs directly with respect to their input parameters. For a detailed discussion of normalizing flows, see Ref. [9089305, 2019arXiv191202762P].
Figure 2: Absolute magnitude $M_{G}$ versus color BPRPBP-RP of all Gaia DR3 stars within 4 kpc of the Sun with full kinematic information, photometric magnitudes, and relative parallax errors $\delta\varpi/\varpi<0.33$. The 4 kpc completeness cut is shown as a white dashed line: stars above this line outside the hatched region are used in our MAF training sample. The Red Clump (RC) – the largest population used in this analysis – is outlined in a white box. Figure reproduced from the authors’ previous work, Ref. [lim2023mapping].
The MAFs used in this work are similar, with minor updates, to those used in our previous work (Ref. [lim2023mapping]). The modifications are:
@@
    Our data preprocessing has been updated to eliminate bias from the autoregressive structure of the MAF. As noted in Ref. [lim2023mapping], MAFs are sensitive to the provided ordering of inputs (in our case, the ordering of Cartesian directions). To counteract this, we inserted Cartesian coordinate permutations between the layers of the MAF. However, since these permutations were identical across training realizations, it is possible that some residual autoregressive bias remained. In this work, we have introduced a further random rotation of the input $\{\vec{x},\vec{v}\}$ of the RC/RGB dataset after centering (${\overline{x}_{i}}^{\prime}=0$) and standardization ($\sigma_{x_{i}}^{\prime}=1$) for each training realization.
@@
Aside from these minor changes, the other details of the MAF models are unchanged, in particular they have the same size and structure as in Ref. [lim2023mapping]. The MAFs for both $n(\vec{x})$ and $p(\vec{v}|\vec{x})$ are composed of two MADE blocks, each with 10 hidden layers that are 48 nodes wide. All MAF models are trained by minimizing the negative log-likelihood (NLL) of the Gaia DR3 RC/RGB dataset introduced in Section 2. For a complete discussion of the MAF training procedure, refer to Ref. [lim2023mapping].
@@ 3.2 Training Dust Efficiency and Potential
We parameterize the functions $\Phi(\vec{x})$ and $\ln\epsilon(\vec{x})$ with fully-connected NNs. Both use five hidden layers, each 100 neurons wide, with the same $\rm GINT$} activations used for the MAFs. An additional transformation layer is included in the $\ln\epsilon$ network to smoothly map the output from $\mathbb{R}$ to $(-\infty,0)$ such that $\epsilon$ is constrained within (0,1)(0,1). [^2] We choose the $C^{\infty}$ function $y=-\frac{1}{\pi}\left(x\left(\arctan{x}+\pi/2\right)+1\right)$ for this purpose. The two networks are trained by minimizing the mean-square error (MSE) of CBE in Eq. (1.3) with respect to the NN parameters $\bm{\theta}$ and $\bm{\vartheta}$: